OpenStudy (emily_98):

The product of two consecutive integers is 342. Which quadratic equation can be used to find x, the greater number? hellpppppppppppppp

OpenStudy (emily_98):

x2 + 1 = 342 x2 − 1 = 342 x2 − x + 342 = 0 x2 − x − 342 = 0

OpenStudy (anonymous):

If the bigger one is x, the smaller one is (x-1). The product of these two are x*(x-1)=x^2-x=342. => x^2 - x - 342= 0